OpenStudy (anonymous):

Madison is financing $382,300 to purchase a house. How much money will she save over the life of a 30-year, fixed-rate loan by buying 3 points with a rate of 5.13% instead of not buying points with a rate of 5.88%? Round to the nearest dollar. A) $60,948 B) $63,242 C) $63,242 D) $53,302 (Not a typo)

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

are you familiar with the concept of points in mortgages?

OpenStudy (anonymous):

Not a clue

jimthompson5910 (jim_thompson5910):

points reduce the interest rate

OpenStudy (anonymous):

ok...

jimthompson5910 (jim_thompson5910):

1 point usually reduces the interest rate by 1/4 of a percent or 1/8 of a percent

jimthompson5910 (jim_thompson5910):

so if you don't buy any points up front, the interest rate is 5.88% if you do buy 3 points, then the interest rate is 5.13%

jimthompson5910 (jim_thompson5910):

each point is 1% of the mortgage amount

jimthompson5910 (jim_thompson5910):

each point costs you 1% of the mortgage amount

OpenStudy (anonymous):

got it! so how do I solve this? in stupid words please :)

jimthompson5910 (jim_thompson5910):

1% of $382,300 = 0.01*382,300 = 3,823

jimthompson5910 (jim_thompson5910):

each point costs $3,823

jimthompson5910 (jim_thompson5910):

but it looks like we're ignoring this cost for now

jimthompson5910 (jim_thompson5910):

what is the monthly payment if... mortgage amount = $382,300 number of years = 30 annual interest rate = 5.88%

OpenStudy (anonymous):

would I use the calculator from the bankrate website for mortgage?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

ok good in that case, first payment I got 2,082.75

jimthompson5910 (jim_thompson5910):

is that with the points? or without?

OpenStudy (anonymous):

it's with the points making the rate 5.13%

jimthompson5910 (jim_thompson5910):

ok how about without the points

OpenStudy (anonymous):

Without the points making the rate 5.88% the payment is 2,262.67

jimthompson5910 (jim_thompson5910):

I'm getting the same monthly payments

OpenStudy (anonymous):

ok. now what happens?

jimthompson5910 (jim_thompson5910):

1 point costs 1% of the mortgage how much does 3 points cost?

OpenStudy (anonymous):

3%?

jimthompson5910 (jim_thompson5910):

take 3% of the mortgage

jimthompson5910 (jim_thompson5910):

3% of $382,300 = ??

OpenStudy (anonymous):

yes and I got 11,469

jimthompson5910 (jim_thompson5910):

With 3 points Total Cost = (monthly payment)*(# of months) + (cost for points) Total Cost = (2,082.75)*(360) + (11,469) Total Cost = 761,259

jimthompson5910 (jim_thompson5910):

With 0 points Total Cost = (2,262.67)*(360) Total Cost = 814,561.20

jimthompson5910 (jim_thompson5910):

Amount you save = (total cost with 0 points) - (total cost with 3 points)

OpenStudy (anonymous):

and now I just subtract the two?

jimthompson5910 (jim_thompson5910):

I'm ignoring the down payment because it's the same for both so they subtract and cancel off

OpenStudy (anonymous):

got it! I might have more later but for now thanks again! :)

OpenStudy (anonymous):

My Final Answer and Correct Answer!! $53,302!

jimthompson5910 (jim_thompson5910):

nice work